From information on a previous question: The mean systolic blood pressure for a population of patients (µ) from a local clinic is 130 with a standard deviation (σ) of 18. What systolic blood pressure marks the bottom 20% of lowest reported levels in the population of patients? Rounded to the nearest tenth.

129.3

111.3

114.9

146.7

0.6
Answer #1

Solution:-

Given that,

mean = = 130

standard deviation = = 18

Using standard normal table,

P(Z < z) = 20 %

= P(Z < z) = 0.20  

= P(Z <- 0.84 ) = 0.20

z = - 0.84

Using z-score formula,

x = z * +

x = - 0.84 * 18 +130

x = 114.9